# Heterogeneous Multi-agent Zero-Shot Coordination by Coevolution

### v2

@@ V-D RQ2: How does MAZE perform compared with recently proposed algorithms?
Table II shows the detailed results, i.e., the mean and standard deviation of the reward achieved by each algorithm under each combination of layout and partner. We compute the rank of each algorithm under each setting as in [10], which are averaged in the last row of Table II. Besides, we apply the Wilcoxon rank-sum test with significance level 0.05, to compare MAZE with other methods. We can observe the order of rank “MAZE $<$ MEP $<$ TrajeDi $<$ FCP $<$ PP $<$ SP”, which is consistent with previous observations, e.g., “FCP $<$ PP $<$ SP” in Strouse et al. [42] and “MEP $<$ TrajeDi $<$ FCP $<$ SP” in Zhao et al. [54]. As expected, PP using a population is better than SP using a single individual. The superiority of TrajeDi, FCP, and MEP over PP discloses the advantage of exposing the agent to diverse partners during the training process. The proposed method MAZE performs the best overall. Furthermore, its superiority on the four heterogeneous layouts (i.e., H-CR, AA, AA-2 and FC), suggesting that MAZE is suitable for heterogeneous ZSC. The generally good performance of MAZE with different partners also shows that MAZE can coordinate with partners with different skill levels. We can also observe that for each algorithm on each layout, the partner trained by MAZE is almost always the best. Besides, the MAZE agent paired with the MAZE partner achieves the highest performance on the four heterogeneous layouts. These observations also show the advantage of considering the heterogeneity by MAZE. When paired with Human Proxy partners, the performance of each algorithm is relatively bad, which may be because the human proxy is trained ignoring the heterogeneity and thus hard to coordinate with.
@@ V-E RQ3: How do different components of MAZE influence its performance?
As discussed in Section IV-A, MAZE employs four pairing strategies, namely fixed, best, worst, and random pairing, to pair an agent with a partner from the partner population. The detailed results of MAZE using these four pairing strategies are presented in Table III. Notably, it can be observed that MAZE achieves the best performance when employing random pairing, whereas fixed pairing yields the worst performance. This outcome is in line with expectations, as pairing with a greater variety of partners during training enhances the ability to effectively coordinate with novel and unseen partners. The design of an improved pairing strategy remains an interesting area for further research.
Table III: The reward (mean±$\pm$±std.) achieved by MAZE using different pairing strategies. For each combination of layout and partner, the largest reward is bolded.
@@
In the updating process of MAZE, we incorporate a diversity term into the objective function to ensure diversity within both the agent and partner populations, as discussed in Section IV-B. By default, we utilize Jensen-Shannon Divergence (JSD) to measure the diversity of the policy population. Additionally, we also employ trajectory-based diversity, as introduced in TrajeDi [26], which calculates diversity based on the trajectory distributions of policies rather than the action distributions used in JSD. To examine the impact of the diversity term on experimental results, we introduce MAZE-T, which incorporates trajectory-based diversity as the diversity term of MAZE, and compare it with TrajeDi and the original MAZE. As shown in Table IV, we observe that MAZE consistently outperforms the other methods, including MAZE-T and TrajeDi. Interestingly, despite both TrajeDi and MAZE-T utilizing the same diversity measure, the use of the MAZE framework significantly improves the performance of MAZE-T in comparison to TrajeDi. Moreover, the performance of MAZE and MAZE-T is similar, suggesting that the diversity term has minimal effect on performance in this context. This finding highlights the need for further investigation into the development of diversity measures that are better suited for addressing the ZSC problem in the future.
Table IV: The reward (mean±$\pm$±std.) achieved by TrajeDi, MAZE with trajectory diversity (MAZE-T) and MAZE. For each combination of layout and partner, the largest reward is bolded.
@@
We first examine the influence of deployment strategies on the performance of MAZE. When terminating, MAZE gets a population of diverse agents rather than a single one. Table V shows the performance of MAZE using two different strategies (i.e., Ensemble and BRO) to deploy the agent population. Given a state, the Ensemble strategy selects the action which receives the most vote from the agents in the population. BRO evaluates each agent in the population using the partners archived by MAZE during training, and selects the best one for testing. We can observe that Ensemble is clearly better than BRO. For each combination of layout and partner, we also test the performance of each agent in the population and report their best and average performance, as shown in the last two columns of Table V. The average rank of Ensemble is very close to that of Best (i.e., the true best single agent, which is not known in advance), showing the robustness of the Ensemble strategy.
Table V: The reward (mean±$\pm$±std.) achieved by MAZE using different deployment strategies. For each combination of layout and partner, the largest reward is bolded.
@@
As we introduce before, MAZE has three main components, i.e., the random pairing of agents and partners (denoted as P), updating each policy by optimizing an objective function with both reward and diversity (denote as D), and partner selection using the archive (denoted as S). To examine the influence of different components of MAZE, we start from the simplest V-MAZE and add these components gradually, generating a series of variants until the complete MAZE, which are denoted as V-MAZE, V-MAZE${}_{+P}$, V-MAZE${}_{+D}$, V-MAZE${}_{+P,D}$, and V-MAZE${}_{+P,D,S}$ (i.e., MAZE). Note that V-MAZE is the simplest variant of MAZE, which uses the fixed pairing strategy, updates the policies without considering the diversity term in the objective function, and does not use the archive for partner selection. The performance comparison among these methods is shown in Table VI. From the last row, i.e., average rank, we can find that using more components generally yields better performance. V-MAZE${}_{+D}$ has a similar performance to V-MAZE, because only adding the diversity term into the objective function without changing pairing cannot ensure that the agents are trained with diverse partners. V-MAZE${}_{+P}$ also adds a single component (i.e., random pairing of agents and partners) compared to V-MAZE, but is better than V-MAZE and V-MAZE${}_{+D}$. V-MAZE${}_{+P}$ changes the partners in pairing, and thus the collected trajectories of each agent-partner pair are more diverse than before, making the agent more likely to coordinate with novel unseen partners well. By combining both the two components, V-MAZE${}_{+P,D}$ performs much better. Finally, the complete MAZE (i.e., V-MAZE${}_{+P,D,S}$) has the best average rank, and the largest reward in each heterogeneous environment (i.e., 183.7 in H-CR, 315.5 in AA, 381.4 in AA-2, and 164.0 in FC) is all achieved by MAZE. Thus, these results indicate the necessity and effectiveness of each component of the proposed framework.
@@
Table VI: The reward (mean±$\pm$±std.) achieved by MAZE using different components. For each combination of layout and partner, the largest reward is bolded.
Layout Partner V-MAZE V-MAZE${}_{+P}$ V-MAZE${}_{+D}$ V-MAZE${}_{+P,D}$ MAZE
H-CR Random 4.0 $\pm$ 0.7 4.4 $\pm$ 1.1 5.8 $\pm$ 0.8 4.5 $\pm$ 0.5 7.0 $\pm$ 0.7
Self Play 50.3 $\pm$ 7.1 86.1 $\pm$ 6.4 55.4 $\pm$ 11.2 101.6 $\pm$ 13.1 97.5 $\pm$ 14.6
MAZE 124.5 $\pm$ 14.2 152.0 $\pm$ 11.4 131.4 $\pm$ 14.6 160.2 $\pm$ 14.5 183.7 $\pm$ 11.9
Human Proxy 69.5 $\pm$ 10.1 94.6 $\pm$ 18.9 72.0 $\pm$ 13.0 101.8 $\pm$ 16.3 118.6 $\pm$ 13.2
AA Random 48.0 $\pm$ 8.7 52.0 $\pm$ 13.5 57.0 $\pm$ 16.1 60.3 $\pm$ 10.9 74.4 $\pm$ 15.5
Self Play 207.0 $\pm$ 20.9 151.7 $\pm$ 36.9 122.0 $\pm$ 23.4 168.5 $\pm$ 35.1 142.6 $\pm$ 20.3
MAZE 310.5 $\pm$ 52.6 286.0 $\pm$ 16.1 262.2 $\pm$ 37.4 311.5 $\pm$ 17.4 315.5 $\pm$ 10.1
Human Proxy 43.5 $\pm$ 6.0 101.3 $\pm$ 17.5 82.5 $\pm$ 22.2 126.3 $\pm$ 27.3 112.0 $\pm$ 19.1
AA-2 Random 52.0 $\pm$ 10.6 59.6 $\pm$ 20.2 49.8 $\pm$ 12.8 53.0 $\pm$ 11.4 56.5 $\pm$ 9.3
Self Play 123.4 $\pm$ 29.9 100.8 $\pm$ 25.4 117.0 $\pm$ 27.8 115.6 $\pm$ 28.9 130.6 $\pm$ 18.5
MAZE 308.2 $\pm$ 34.3 362.5 $\pm$ 29.2 327.3 $\pm$ 36.4 343.0 $\pm$ 29.5 381.4 $\pm$ 13.7
Human Proxy 60.4 $\pm$ 14.3 75.3 $\pm$ 19.8 80.8 $\pm$ 18.0 104.6 $\pm$ 12.5 111.5 $\pm$ 13.7
FC Random 4.0 $\pm$ 1.9 2.7 $\pm$ 1.1 3.0 $\pm$ 2.0 4.0 $\pm$ 0.4 7.3 $\pm$ 0.4
Self Play 99.0 $\pm$ 37.9 123.0 $\pm$ 34.1 79.4 $\pm$ 24.0 113.3 $\pm$ 35.5 147.0 $\pm$ 27.1
MAZE 108.5 $\pm$ 26.0 131.8 $\pm$ 21.4 93.4 $\pm$ 20.0 123.0 $\pm$ 29.1 164.0 $\pm$ 27.0
Human Proxy 18.3 $\pm$ 5.2 11.3 $\pm$ 4.7 17.8 $\pm$ 2.6 16.3 $\pm$ 4.7 25.0 $\pm$ 5.1
Average Rank 3.88 3.31 3.94 2.50 1.38
@@ V-G Further Studies
Recent work shows that exposing the agent to diverse partners during the training process efficiently improves the ability to coordinate with unseen partners [26, 42, 54]. To verify if the obtained partners are diverse, we visualize the population of partners obtained by different methods after the training phase. For a fair comparison, we pair each partner with the same agent and record the partner’s action trajectory. We repeat this five times and then visualize all the obtained trajectories in a two-dimensional space using t-SNE [45]. The result is shown in Figure 9. Note that there are 25 trajectories (i.e., 25 points) for the population-based methods (five partners obtained by each method and five trajectories for each partner). The partners obtained by MAZE are the most diverse because they are distributed in different areas of the embedded space. The partners obtained by TrajeDi, FCP, and MEP are also distributed to some extent but not as diverse as MAZE. This is because their mechanisms for maintaining partner diversity are less effective than MAZE. As a population-based method, PP’s partners are not diverse enough due to the lack of explicit mechanisms to maintain diversity. The partners obtained by SP are concentrated, which is consistent with its worst performance in RQ1.
@@
Figure 10: Training curves of SP, PP and V-MAZE on CR and H-CR.

### v3

@@ V-D RQ2: How does MAZE perform compared with recently proposed algorithms?
Table III shows the detailed results, i.e., the mean and standard deviation of the reward achieved by each algorithm under each combination of layout and partner. We compute the rank of each algorithm under each setting as in [10], which are averaged in the last row of Table III. Besides, we apply the Wilcoxon rank-sum test with significance level 0.05, to compare MAZE with other methods. We can observe the order of rank “MAZE $<$ MEP $<$ TrajeDi $<$ FCP $<$ PP $<$ SP”, which is consistent with previous observations, e.g., “FCP $<$ PP $<$ SP” in Strouse et al. [48] and “MEP $<$ TrajeDi $<$ FCP $<$ SP” in Zhao et al. [62]. As expected, PP using a population is better than SP using a single individual. The superiority of TrajeDi, FCP, and MEP over PP discloses the advantage of exposing the agent to diverse partners during the training process. The proposed method MAZE performs the best overall. Furthermore, its superiority on the four heterogeneous layouts (i.e., H-CR, AA, AA-2 and FC), suggesting that MAZE is suitable for heterogeneous ZSC. The generally good performance of MAZE with different partners also shows that MAZE can coordinate with partners with different skill levels. We can also observe that for each algorithm on each layout, the partner trained by MAZE is almost always the best. Besides, the MAZE agent paired with the MAZE partner achieves the highest performance on the four heterogeneous layouts. These observations also show the advantage of considering the heterogeneity by MAZE. When paired with Human Proxy partners, the performance of each algorithm is relatively bad, which may be because the human proxy is trained ignoring the heterogeneity and thus hard to coordinate with.
@@ V-E RQ3: How do different components of MAZE influence its performance?
As discussed in Section IV-A, MAZE employs four pairing strategies, namely fixed, best, worst, and random pairing, to pair an agent with a partner from the partner population. The detailed results of MAZE using these four pairing strategies are presented in Table IV. Notably, it can be observed that MAZE achieves the best performance when employing random pairing, whereas fixed pairing yields the worst performance. This outcome is in line with expectations, as pairing with a greater variety of partners during training enhances the ability to effectively coordinate with novel and unseen partners. The design of an improved pairing strategy remains an interesting area for further research.
Table IV: The reward (mean±$\pm$±std.) achieved by MAZE using different pairing strategies. For each combination of layout and partner, the largest reward is bolded.
@@
In the updating process of MAZE, we incorporate a diversity term into the objective function to ensure diversity within both the agent and partner populations, as discussed in Section IV-B. By default, we utilize Jensen-Shannon Divergence (JSD) to measure the diversity of the policy population. Additionally, we also employ trajectory-based diversity, as introduced in TrajeDi [31], which calculates diversity based on the trajectory distributions of policies rather than the action distributions used in JSD. To examine the impact of the diversity term on experimental results, we introduce MAZE-T, which incorporates trajectory-based diversity as the diversity term of MAZE, and compare it with TrajeDi and the original MAZE. As shown in Table V, we observe that MAZE consistently outperforms the other methods, including MAZE-T and TrajeDi. Interestingly, despite both TrajeDi and MAZE-T utilizing the same diversity measure, the use of the MAZE framework significantly improves the performance of MAZE-T in comparison to TrajeDi. Moreover, the performance of MAZE and MAZE-T is similar, suggesting that the diversity term has minimal effect on performance in this context. This finding highlights the need for further investigation into the development of diversity measures that are better suited for addressing the ZSC problem in the future.
Table V: The reward (mean±$\pm$±std.) achieved by TrajeDi, MAZE with trajectory diversity (MAZE-T) and MAZE. For each combination of layout and partner, the largest reward is bolded.
@@
We first examine the influence of deployment strategies on the performance of MAZE. When terminating, MAZE gets a population of diverse agents rather than a single one. Table VI shows the performance of MAZE using two different strategies (i.e., Ensemble and BRO) to deploy the agent population. Given a state, the Ensemble strategy selects the action which receives the most vote from the agents in the population. BRO evaluates each agent in the population using the partners archived by MAZE during training, and selects the best one for testing. We can observe that Ensemble is clearly better than BRO. For each combination of layout and partner, we also test the performance of each agent in the population and report their best and average performance, as shown in the last two columns of Table VI. The average rank of Ensemble is very close to that of Best (i.e., the true best single agent, which is not known in advance), showing the robustness of the Ensemble strategy.
Table VI: The reward (mean±$\pm$±std.) achieved by MAZE using different deployment strategies. For each combination of layout and partner, the largest reward is bolded.
@@
As we introduce before, MAZE has three main components, i.e., the random pairing of agents and partners (denoted as P), updating each policy by optimizing an objective function with both reward and diversity (denote as D), and partner selection using the archive (denoted as S). To examine the influence of different components of MAZE, we start from the simplest V-MAZE and add these components gradually, generating a series of variants until the complete MAZE, which are denoted as V-MAZE, V-MAZE+P, V-MAZE+D, V-MAZE+S, V-MAZE+S,P, V-MAZE+S,D, V-MAZE+P,D, and V-MAZE+P,D,S (i.e., MAZE). Note that V-MAZE is the simplest variant of MAZE, which uses the fixed pairing strategy, updates the policies without considering the diversity term in the objective function, and does not use the archive for partner selection. The performance comparison among these methods is shown in Table VII. From the last row, i.e., average rank, we can find that using more components generally yields better performance. V-MAZE+D has a similar performance to V-MAZE, because only adding the diversity term into the objective function without changing pairing cannot ensure that the agents are trained with diverse partners. V-MAZE+P also adds a single component (i.e., random pairing of agents and partners) compared to V-MAZE, but is better than V-MAZE and V-MAZE+D. V-MAZE+P changes the partners in pairing, and thus the collected trajectories of each agent-partner pair are more diverse than before, making the agent more likely to coordinate with novel unseen partners well. By combining both the two components, V-MAZE+P,D performs much better. In V-MAZE+S, partners are chosen from the archive each generation, allowing the agent to encounter different partners and improve final coordination ability. V-MAZE+S,P combines selection and pairing, further promoting the agent’s exposure to a diverse range of partners. V-MAZE+S,D additionally uses diversity-based updates, making the partners even more diverse and further enhancing the final results. Finally, the complete MAZE (i.e., V-MAZE+P,D,S) has the best average rank, and the largest reward in each heterogeneous environment (i.e., 183.7 in H-CR, 315.5 in AA, 381.4 in AA-2, and 164.0 in FC) is all achieved by MAZE. Thus, these results indicate the necessity and effectiveness of each component of the proposed framework.
@@
Table VII: The reward (mean±$\pm$±std.) achieved by MAZE using different components. For each combination of layout and partner, the largest reward is bolded.
Layout Partner V-MAZE V-MAZE+P V-MAZE+D V-MAZE+S V-MAZE+S,P V-MAZE+S,D V-MAZE+P,D MAZE
H-CR Random 4.0 $\pm$ 0.7 4.4 $\pm$ 1.1 5.8 $\pm$ 0.8 5.4 $\pm$ 0.6 5.7 $\pm$ 0.5 4.9 $\pm$ 0.6 4.5 $\pm$ 0.5 7.0 $\pm$ 0.7
Self Play 50.3 $\pm$ 7.1 86.1 $\pm$ 6.4 55.4 $\pm$ 11.2 54.5 $\pm$ 10.1 58.9 $\pm$ 9.2 87.3 $\pm$ 12.4 101.6 $\pm$ 13.1 97.5 $\pm$ 14.6
MAZE 124.5 $\pm$ 14.2 152.0 $\pm$ 11.4 131.4 $\pm$ 14.6 129.7 $\pm$ 12.3 145.1 $\pm$ 13.5 156.2 $\pm$ 15.1 160.2 $\pm$ 14.5 183.7 $\pm$ 11.9
Human Proxy 69.5 $\pm$ 10.1 94.6 $\pm$ 18.9 72.0 $\pm$ 13.0 86.1 $\pm$ 11.5 76.2 $\pm$ 12.4 103.4 $\pm$ 14.8 101.8 $\pm$ 16.3 118.6 $\pm$ 13.2
AA Random 48.0 $\pm$ 8.7 52.0 $\pm$ 13.5 57.0 $\pm$ 16.1 55.4 $\pm$ 9.3 57.3 $\pm$ 12.1 61.2 $\pm$ 11.7 60.3 $\pm$ 10.9 74.4 $\pm$ 15.5
Self Play 207.0 $\pm$ 20.9 151.7 $\pm$ 36.9 122.0 $\pm$ 23.4 178.1 $\pm$ 33.4 145.1 $\pm$ 31.2 182.3 $\pm$ 27.9 168.5 $\pm$ 35.1 142.6 $\pm$ 20.3
MAZE 310.5 $\pm$ 52.6 286.0 $\pm$ 16.1 262.2 $\pm$ 37.4 251.7 $\pm$ 32.1 249.8 $\pm$ 30.0 301.2 $\pm$ 21.4 311.5 $\pm$ 17.4 315.5 $\pm$ 10.1
Human Proxy 43.5 $\pm$ 6.0 101.3 $\pm$ 17.5 82.5 $\pm$ 22.2 113.5 $\pm$ 21.7 117.2 $\pm$ 17.2 122.3 $\pm$ 16.7 126.3 $\pm$ 27.3 112.0 $\pm$ 19.1
AA-2 Random 52.0 $\pm$ 10.6 59.6 $\pm$ 20.2 49.8 $\pm$ 12.8 49.1 $\pm$ 11.3 62.3 $\pm$ 21.1 59.3 $\pm$ 19.4 53.0 $\pm$ 11.4 56.5 $\pm$ 9.3
Self Play 123.4 $\pm$ 29.9 100.8 $\pm$ 25.4 117.0 $\pm$ 27.8 106.3 $\pm$ 18.9 122.5 $\pm$ 26.7 129.3 $\pm$ 17.7 127.5 $\pm$ 27.3 130.6 $\pm$ 18.5
MAZE 308.2 $\pm$ 34.3 362.5 $\pm$ 29.2 327.3 $\pm$ 36.4 328.2 $\pm$ 27.6 345.7 $\pm$ 26.8 352.3 $\pm$ 31.7 343.0 $\pm$ 29.5 381.4 $\pm$ 13.7
Human Proxy 60.4 $\pm$ 14.3 75.3 $\pm$ 19.8 80.8 $\pm$ 18.0 81.4 $\pm$ 19.4 87.7 $\pm$ 15.1 103.7 $\pm$ 16.3 104.6 $\pm$ 12.5 111.5 $\pm$ 13.7
FC Random 4.0 $\pm$ 1.9 2.7 $\pm$ 1.1 3.0 $\pm$ 2.0 2.3 $\pm$ 1.2 3.6 $\pm$ 1.7 4.5 $\pm$ 0.6 4.0 $\pm$ 0.4 7.3 $\pm$ 0.4
Self Play 99.0 $\pm$ 37.9 123.0 $\pm$ 34.1 79.4 $\pm$ 24.0 82.1 $\pm$ 34.7 121.3 $\pm$ 27.2 132.1 $\pm$ 29.4 113.3 $\pm$ 35.5 147.0 $\pm$ 27.1
MAZE 108.5 $\pm$ 26.0 131.8 $\pm$ 21.4 93.4 $\pm$ 20.0 112.5 $\pm$ 18.7 114.3 $\pm$ 17.9 119.1 $\pm$ 28.2 123.0 $\pm$ 29.1 164.0 $\pm$ 27.0
Human Proxy 18.3 $\pm$ 5.2 11.3 $\pm$ 4.7 17.8 $\pm$ 2.6 15.4 $\pm$ 3.2 20.3 $\pm$ 4.9 17.9 $\pm$ 5.1 16.3 $\pm$ 4.7 25.0 $\pm$ 5.1
Average Rank 6.13 5.06 6.25 6.06 4.37 2.88 3.37 1.87
Sensitivity analysis.
MAZE has two hyper-parameters: population size $n_{P}$ and and archive size $n_{A}$, which is set to $5$ and $20$ by default. We further conduct experiments to analysis their influence on the final performance. We compare the default MAZE with MAZE-pop3 (i.e., the population size is set to 3, with the archive size remaining unchanged at 20), MAZE-pop7, MAZE-archive10 (i.e., the archive size is set to 10, with the population size remaining unchanged at 5), and MAZE-archive40. The results can be found in Table VIII. For a fair experimental comparison, we keep the number of updates for each agent consistent. Therefore, changing the population size will result in different total number of updates, which may allow larger populations to achieve relatively better results (i.e., the more agents we have, the easier it is to find better-performing agents). When the archive size increases, the number of updates for each potential partner decreases, leading to slightly lower performance in the same number of iterations. However, compared to other comparison algorithms in the main experiments (i.e., Table III), all the MAZE variants still maintain good performance.
Table VIII: The reward (mean±$\pm$±std.) achieved by MAZE with different population size and archive size. MAZE uses population size 5 and archive size 20 by default. MAZE-pop3 and MAZE-pop7 denote MAZE with population size 3 and 7, respectively; MAZE-archive10 and MAZE-archive40 denote MAZE with archive size 10 and 40, respectively. For each combination of layout and partner, the largest reward is bolded.
Layout Partner MAZE MAZE-pop3 MAZE-pop7 MAZE-archive10 MAZE-archive40
AA-2 Random 56.5 $\pm$ 9.3 59.3 $\pm$ 8.5 49.2 $\pm$ 7.3 55.2 $\pm$ 7.1 53.4 $\pm$ 9.1
Self-Play 130.6 $\pm$ 18.5 127.2 $\pm$ 14.7 121.2 $\pm$ 17.9 124.5 $\pm$ 14.7 119.3 $\pm$ 14.6
MAZE 381.4 $\pm$ 13.7 382.8 $\pm$ 11.2 389.1 $\pm$ 14.1 371.0 $\pm$ 12.5 376.4 $\pm$ 8.6
Human Proxy 111.5 $\pm$ 13.7 109.4 $\pm$ 16.8 98.4 $\pm$ 7.3 113.2 $\pm$ 12.9 87.2 $\pm$ 9.6
FC Random 7.3 $\pm$ 0.4 5.2 $\pm$ 0.1 7.7 $\pm$ 0.2 7.4 $\pm$ 0.7 6.1 $\pm$ 0.9
Self-Play 147.0 $\pm$ 27.1 138.1 $\pm$ 26.3 146.1 $\pm$ 20.4 141.2 $\pm$ 19.8 132.0 $\pm$ 21.5
MAZE 164.0 $\pm$ 27.0 153.2 $\pm$ 25.4 161.2 $\pm$ 26.1 165.1 $\pm$ 27.7 156.9 $\pm$ 22.1
Human Proxy 25.0 $\pm$ 5.1 19.2 $\pm$ 4.8 27.1 $\pm$ 6.3 26.1 $\pm$ 4.2 20.4 $\pm$ 6.3
Average Rank 2.13 3.37 2.63 2.5 4.37
@@ V-G Further Studies
Recent work shows that exposing the agent to diverse partners during the training process efficiently improves the ability to coordinate with unseen partners [31, 48, 62]. To verify if the obtained partners are diverse, we visualize the population of partners obtained by different methods after the training phase. For a fair comparison, we pair each partner with the same agent and record the partner’s action trajectory. We repeat this five times and then visualize all the obtained trajectories in a two-dimensional space using t-SNE [51]. The result is shown in Figure 9. Note that there are 25 trajectories (i.e., 25 points) for the population-based methods (five partners obtained by each method and five trajectories for each partner). Some points are clustered nearby in the figure, because we repeated the experiments with same agents multiple times to obtain multiple training trajectories. The partners obtained by MAZE are the most diverse because they are distributed in different areas of the embedded space. xxThe partners obtained by TrajeDi, FCP, and MEP are also distributed to some extent but not as diverse as MAZE. This is because their mechanisms for maintaining partner diversity are less effective than MAZE. As a population-based method, PP’s partners are not diverse enough due to the lack of explicit mechanisms to maintain diversity. The partners obtained by SP are concentrated, which is consistent with its worst performance in RQ1.
@@
Figure 10: Training curves on CR and H-CR.